Observer: Fuyan Bian, Cai Zheng, Ran Wang. Condition: Cloudy in the first half night.

08:37 we open the telescope, go to zheng's targets mask1_PA50, something wrong with the mask alignment, go to mask2_PA50, everything is fine. Seeing is 1.0-1.2 at the beginning of exposure, then figured out the first problem on mask alignment was due to an operating problem. lose the guide star due to cloud and worse seeing after 20 minutes exposure. have to stop observing, finish one 20-min exposure.

10:20 go to SN2001em, seeing is 1.5 and passing thin clouds, finished 40-min exposure.

11:30 go to bd28 standard star.

-- FuyanBian - 15 May 2013
Topic revision: r3 - 15 May 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback